OpenStudy (anonymous):

Need serious help with orthogonal projections?? A linearly dependent set of vectors in R^4 is given. Find the standard matrix for the orthogonal projection of R^4 onto the subspace spanned by those vectors?

OpenStudy (anonymous):

where a_1=(4, -6, 2 ,-4) a_2=(2,-3, 1,-2) a_3=(1,0,-2,5) a_4=(5,-6,0,1)

OpenStudy (anonymous):

to me it sounds like you want to create an orthogonal basis from those vectors give. This requires the Gram-Schmit process. Do you know about that, or are you familiar with it?

OpenStudy (anonymous):

Well I know what that is but that's the chapter after this so I don't think we use Gram Schmit process

OpenStudy (anonymous):

if its the chapter after it should be fair game no? I could understand if it was the chapter before....its like you havent learned it yet. But after, it should be considered a tool for you to use.

OpenStudy (anonymous):

No haha I meant this question is the chapter before learning Gram Schmit, but yeah is that the quickest way then, I'm just studying for my final so whichever way is the quickest would be best

OpenStudy (anonymous):

im not too sure how you would do it without gram-schmit....i'll think about it for a bit lol. but yeah that would be the most straightforward way to tackle this problem.

OpenStudy (anonymous):

Like if it was only one vector there's this formula (1/(A^T*A))*(A*A^T)

OpenStudy (anonymous):

and there is also this formula M(M^TM)^-1*M^T where M is any matrix. So I thought I could combine the vectors to make a matrix but not getting the answer

OpenStudy (anonymous):

Oh nvm I got it you do use the M formula, I put the vectors in rows but it's actually columns so it works now.

OpenStudy (anonymous):

hmmm, that sounds interesting, im gonna take a look at it too now, nice job!